# Sensing for Space Safety and Sustainability: A Deep Learning Approach with Vision Transformers

### v1

@@ I Introduction
With the rapid increase of advanced low-Earth-orbit (LEO) satellites in large constellations, space assets are set to play a crucial role in the future, enabling global Internet access and space relay systems for deep space missions. However, managing thousands of LEO satellites with the assurance of safety operations, space sustainability, and space situational awareness is becoming a significant challenge. For example, the increasing risk of collisions between LEO satellites and space objects can cause significant amount of space debris in various sizes and threaten the safety operations of spacecraft and outer space environment. Theses risks are real and will be exacerbated with the increasing space assets into the outer space. As evident from the past incidents between satellites and space debris and from the study by the Inter-Agency Space Debris Coordination Committee (IADC) [1] indicating that “catastrophic collisions are expected every 5 to 9 years” in an almost perfect scenario.
@@ III System Model
(a) The architecture of the GELAN-ViT model.
(b) The architecture of the GELAN-RepViT model.
Figure 1: Architectures of the proposed GELAN-ViT and GELAN-RepViT models.
@@ III-A Overview
Our work focuses on enhancing the performance of GELAN-t by addressing the inefficiencies caused by combining local features ($F_{l}$) and global features ($F_{g}$)) within the same pathway. Previous work has shown that when $F_{l}$ and $F_{g}$ are processed together, they compete for limited neural resources, decreasing performance. This issue arises from the information bottleneck principle, which states that neural networks have a finite capacity for transmitting information [12]. When both feature types are forced to share the same pathway, the network’s capacity must be divided between them, limiting the amount of information that can be retained for both local and global features as shown in (1):
@@ III-B2 GELAN-RepViT
The GELAN-RepViT architecture, shown in Fig. 1(b), is an extension of the GELAN-s model with three modifications. First, similar to GELAN-ViT, the number of neurons in each layer is scaled down by a factor of 0.25. Second, the architecture is simplified by removing a redundant AnyBlock module to reduce complexity. Third, we add a dedicated branch that integrates RepNCSPELAN4_ViT, a version of RepNCSPELAN4 integrated with ViT encoder layers, as proposed by [6].
@@ IV Experiments
In the experiments, we compared our proposed model with GELAN-t [4], YOLOv9-t [4], YOLOv7-t [3], YOLOX-s [2], YOLOX-ViT-s [6], and MobileViT [5] using the SOD and VOC 2012 datasets. We utilized an HPC cluster with an NVIDIA V100 16 GB for training and testing. The results are averaged over 20 test runs with an image size of 640 x 640 for the SOD dataset and 320 x 320 for the VOC 2012 dataset. The current SOD dataset is designed for SOD tasks for LEO satellites and consists of satellite object classes and celestial bodies. In contrast, the VOC 2012 dataset is a commonly used benchmark dataset for object detection tasks, which contains 20 different object classes. The VOC 2012 dataset is more complex due to the variability of the classes.
@@ IV-A Evaluation Metrics
Our evaluation strategy focuses on key metrics in object detection tasks with regard to Intersection over Union (IoU), specifically,mAP50mAP50:95. The metrics we use are defined below.
  • Average Precision ($AP_{i}$): This is calculated as the area under the precision-recall curve for the IoU threshold $i$ ($p_{i}(r)$).
@@ IV-C Performance on SOD Dataset
The performance of our models in the SOD dataset (SODD) is summarized in Table I. Our proposed models, GELAN-ViT and GELAN-RepViT, are compared with GELAN-t, YOLOv9-t, YOLOv7-t, YOLOX-s, YOLOX-ViT-s, and MobileViT. The metrics used for the evaluation include mAP50 at 1000 epochs, mAP50:95 at 1000 epochs, GFLOPs, and the number of parameters. In calculating GFLOPs, YOLOv9 and YOLOv9-t default to an input size of 640 x 640, while MobileViT, YOLOX, and YOLOX-ViT use their respective input sizes. To ensure consistency in our comparison, we fixed the GFLOPs calculation to an input size of 640 x 640 for all models.
@@ IV-D Performance on VOC 2012 Dataset
The performance of our models in the VOC 2012 dataset is also summarized in Table I. The metrics used for evaluation include mAP50 at 1000 epochs, mAP50:95 at 1000 epochs, GFLOPs, and the number of parameters. The input size for the VOC dataset is standardized to 320 x 320. However, for consistency in GFLOPs calculation across all models, we fixed the GFLOPs calculation to an input size of 640 x 640, regardless of the original input size used in training.

### v2

@@ I Introduction
With the rapid increase of advanced low-Earth-orbit (LEO) satellites in large constellations, space assets are set to play a crucial role in the future, enabling global Internet access and space relay systems for deep space missions. However, managing thousands of LEO satellites with the assurance of safety operations, space sustainability, and space situational awareness is becoming a significant challenge. For example, the increasing risk of collisions between LEO satellites and space objects can cause significant amount of space debris in various sizes and threaten the safety operations of spacecraft and outer space environment. These risks are real and will be exacerbated with the increasing space assets into the outer space. As evident from the past incidents between satellites and space debris and from the study by the Inter-Agency Space Debris Coordination Committee (IADC) [1] indicating that “catastrophic collisions are expected every 5 to 9 years” in an almost perfect scenario.
@@ III System Model
(a) The architecture of the GELAN-ViT model
(b) The architecture of the GELAN-RepViT model
Figure 1: Architectures of the proposed GELAN-ViT and GELAN-RepViT models
@@ III-A Overview
Our work focuses on enhancing the performance of GELAN-t by addressing the inefficiencies caused by combining local features ($F_{l}$) and global features ($F_{g}$) within the same pathway. Previous work has shown that when $F_{l}$ and $F_{g}$ are processed together, they compete for limited neural resources, decreasing performance. This issue arises from the information bottleneck principle, which states that neural networks have a finite capacity for transmitting information [12]. When both feature types are forced to share the same pathway, the network’s capacity must be divided between them, limiting the amount of information that can be retained for both local and global features as shown in (1):
@@ III-B2 GELAN-RepViT
The GELAN-RepViT architecture, shown in Fig. 1(b), is an extension of the GELAN-s model with three modifications. First, similar to GELAN-ViT, the number of neurons in each layer is scaled down by a factor of 0.25. Second, the architecture is simplified by removing a redundant detection head to reduce complexity. Third, we add a dedicated branch that integrates RepNCSPELAN4_ViT, a version of RepNCSPELAN4 integrated with ViT encoder layers, as proposed by [6].
@@ IV Experiments
In the experiments, we compared our proposed model with GELAN-t [4], YOLOv9-t [4], YOLOv7-t [3], YOLOX-s [2], YOLOX-ViT-s [6], and MobileViT [5] using the SOD and VOC 2012 datasets. We utilized an HPC cluster with an NVIDIA V100 16 GB for training and testing. The results are averaged over 20 test runs with an image size of 640 $\times$ 640 for the SOD dataset and 320 $\times$ 320 for the VOC 2012 dataset. The current SOD dataset is designed for SOD tasks for LEO satellites and consists of satellite object classes and celestial bodies. In contrast, the VOC 2012 dataset is a commonly used benchmark dataset for object detection tasks, which contains 20 different object classes. The VOC 2012 dataset is more complex due to the variability of the classes.
@@ IV-A Evaluation Metrics
Our evaluation strategy focuses on key metrics in object detection tasks with regard to Intersection over Union (IoU), specifically, mAP50 and mAP50:95. The metrics we use are defined below.
  • Average Precision ($AP_{u}$): This is calculated as the area under the precision-recall curve for the IoU threshold $u$ ($p_{u}(r)$).
@@ IV-C Performance on SOD Dataset
The performance of our models in the SOD dataset (SODD) is summarized in Table I. Our proposed models, GELAN-ViT and GELAN-RepViT, are compared with GELAN-t, YOLOv9-t, YOLOv7-t, YOLOX-s, YOLOX-ViT-s, and MobileViT. The metrics used for the evaluation include mAP50 at 1000 epochs, mAP50:95 at 1000 epochs, GFLOPs, and the number of parameters. In calculating GFLOPs, YOLOv9 and YOLOv9-t default to an input size of 640 $\times$ 640, while MobileViT, YOLOX, and YOLOX-ViT use their respective input sizes. To ensure consistency in our comparison, we fixed the GFLOPs calculation to an input size of 640 $\times$ 640 for all models.
@@ IV-D Performance on VOC 2012 Dataset
The performance of our models in the VOC 2012 dataset is also summarized in Table I. The metrics used for evaluation include mAP50 at 1000 epochs, mAP50:95 at 1000 epochs, GFLOPs, and the number of parameters. The input size for the VOC dataset is standardized to 320 $\times$ 320. However, for consistency in GFLOPs calculation across all models, we fixed the GFLOPs calculation to an input size of 640 $\times$ 640, regardless of the original input size used in training.